OpenStudy (anonymous):

Find the approximate value of the volume of the right circular cone with a circular base . Approximate your solution to the nearest hundredth. 12cm 10 cm

OpenStudy (anonymous):

what are 12cm and 10cm? volume of a cone = 1/3 times area of base times height.

OpenStudy (anonymous):

V=(1/3)π*r^2*h V=(1/3)π*(1.5)^2*(7) V=16.4934. The cone can hold 16.49 cubic inches of ice cream is for the last

OpenStudy (anonymous):

where'd you get that r=1.5 and h=7 from? your question listed 12cm and 10cm.

OpenStudy (anonymous):

no i didnt say that im tlk about for the last problem how is that the same method

OpenStudy (anonymous):

OK, yes, it's the same method. although you don't say what the 12cm and 10cm are in your question.

OpenStudy (anonymous):

12cm is the bottom of the cone and 10 is the right side

OpenStudy (anonymous):

so diameter =12 and height=10? or is the 10 measured parallel to the side? you need to describe these with enough detail, or else we (you) might get the answers wrong.